OpenStudy (anonymous):

Find the complex zeros of the polynomial functions. Write f in factored form a. f(x)=x^3-216 b. f(x)=x^3-4x^2+14x-20 c. x^4+6x^2+5 **Use the complex zeros to write f in factored form

OpenStudy (anonymous):

@jim_thompson5910 Could you help me with this problem please?

jimthompson5910 (jim_thompson5910):

x^3-216 can be factored using a difference of cubes x^3-216 = x^3 - 6^3 = (x-6)(x^2 + 6*x + 6^2) = (x-6)(x^2 + 6x + 36) the others will have to use a graphing calculator

jimthompson5910 (jim_thompson5910):

actually with part c), you can let z = x^2 which means z^2 = x^4

OpenStudy (anonymous):

it says (x-6)(x^2+6x+36) is wrong

jimthompson5910 (jim_thompson5910):

oh you have to solve x^2+6x+36 = 0 now

OpenStudy (anonymous):

oh okay gotcha

jimthompson5910 (jim_thompson5910):

to get the complex roots

OpenStudy (anonymous):

what am i supposed to get for the complex roots,cause it says im wrong again

jimthompson5910 (jim_thompson5910):

what did you get when you solved x^2+6x+36 = 0

jimthompson5910 (jim_thompson5910):

btw this problem http://assets.openstudy.com/updates/attachments/5459ba7ee4b07e447429d065-mhm120-1415167239637-4.69.png is different from what you originally posted

OpenStudy (anonymous):

yeah i know sorry i accidentally posted that one

OpenStudy (anonymous):

and i got 3(-1-i\[3(-1-\sqrt{3}) and 3(-1+i \sqrt{3})\]

OpenStudy (anonymous):

sorry ignore the first part the "3(-1-i" part

jimthompson5910 (jim_thompson5910):

so x^2+6x+36 factors to \[\Large (x-3(-1-i\sqrt{3}))(x-3(-1+i\sqrt{3}))\] it might be easier to write it like this \[\Large (x+3+3i\sqrt{3})(x+3-3i\sqrt{3})\] this is just for the x^2+6x+36 portion though

OpenStudy (anonymous):

so it's be (x-6) (x+3+3i√3)(x+3−3i√3) ??

OpenStudy (anonymous):

idk what happened there one second

jimthompson5910 (jim_thompson5910):

yeah correct

OpenStudy (anonymous):

\[(x-6)(x+3+3i \sqrt{3})(x+3-3i \sqrt{3})\]

OpenStudy (anonymous):

oh okay thanks!!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

how do you do part c again , i don't understand the z^2=x^4

jimthompson5910 (jim_thompson5910):

Let z = x^2 square both sides to get z^2 = x^4 ------------------------------------------------------- the expression x^4+6x^2+5 turns into z^2+6z+5 after you do substitutions

jimthompson5910 (jim_thompson5910):

now factor z^2+6z+5 to get ???

OpenStudy (anonymous):

(z+3)(z+2)

jimthompson5910 (jim_thompson5910):

now that you know (z+3)(z+2), you would replace each z with x^2 (z+3)(z+2) (x^2+3)(x^2+2)

jimthompson5910 (jim_thompson5910):

you can factor each further, but you'll get complex factors

OpenStudy (anonymous):

o i would get \[(-i \sqrt{3}) and (i \sqrt{3})\]

jimthompson5910 (jim_thompson5910):

for x^2+3, yes

jimthompson5910 (jim_thompson5910):

that means x^2+3 factors to \[\Large (x - i\sqrt{3})(x + i\sqrt{3})\]